OpenStudy (anonymous):

(1- tanQ )(1+ sin2Q)=1+ tanQ

OpenStudy (anonymous):

\[(1-\tan \theta)(1+\sin \theta)=1+\tan \theta\]

sam (.sam.):

prove?

OpenStudy (anonymous):

Yea!

OpenStudy (anonymous):

It's actually sin 2 theta